10. A square has a 40-cm diagonal. How long is each side of the square? Round your answer to the nearest tenth of a centimeter. (1 point
Found 2 solutions by anomis, mananth:
Answer by anomis(18)

All sides of a square are equal, so the diagonal of a square splits the square in two right triangles with two sides that are equal.
We use the Pythagorean Theorem to find the length of the side:
We know that the two sides of the triangle are equal ( because that sides are actually the sides of the square)
So we have:

a^2 + a^2 = 40^2 ( the diagonal of the square is the hypotenuses)
2a^2 = 40^2
a^2 = 800
a = √800
a ~ 28.3 cm
All sides of a square are equal so all sides are 28.3cm

Answer by mananth(16946)

The legs and the diagonal form a right angle.
The diagonal is the hypotenuse.
apply Pythagoras theorem
..
Let l be one leg.
since it is a square the other leg is also l.
l^2+l^2= 40^2
2l^2=1600
l^2=800
l = sqrt (800)
l = 28.28
...
l= 28.3 cm